# COMPASS: Computational Mapping of Patient-Therapist Alliance Strategies with Language Modeling

### v1

@@ 2.6 Analyzing the Temporal Dynamics of Topics
To analyze the temporal dynamics of topics, we compute topic scores at the turn-level. We utilize the Embedded Topic Model (ETM) for this analysis, as it models each word with a categorical distribution based on the inner product between a word embedding and the embedding of its assigned topic [27]. We use the same Word2Vec word embeddings to embed both the topics and the dialogue turns, to then compute the cosine similarity between the embedded topic vector and the embedded turn vector. By applying this methods, which we term Temporal Topic Modeling (TMM), we obtain turn-resolution topic scores that capture the temporal dynamics of the topics discussed during the therapy session. These turn-level topic scores allow us to track the changes in topic relevance over time, providing insights into the progression of the therapy, the emergence of specific topics, and shifts in the focus of the conversation.
@@ 3 Results
3.1 Psychotherapy Transcript Dataset
We begin by introducing the dataset used in our study. The Alex Street Counseling and Psychotherapy Transcripts dataset [16] consists of transcribed recordings of over 950 therapy sessions between multiple anonymized therapists and patients. This comprehensive collection includes speech-translated transcripts of the recordings from real therapy sessions, 40,000 pages of client narratives, and 25,000 pages of reference works. The sessions cover four types of psychiatric conditions: anxiety, depression, schizophrenia, and suicidal. Each dialogue pair consists of a patient response turn followed by a therapist response turn . In total, the dataset contains over 200,000 turns from both patients and therapists, providing a rich source for analyzing the therapeutic process in psychotherapy.
Patient-Therapist Consistency of Working Alliance. We investigate the consistency of the working alliance estimation between patients and therapists. Comparing the estimates, we observe that therapists tend to overestimate the working alliance overall. Specifically, therapists tend to overestimate the task and bond scales, but underestimate the goal scale. These differences are statistically significant (). We also find that the working alliance scores differ significantly between certain pairs of psychiatric conditions, such as anxiety and depression, and anxiety and schizophrenia, in both the therapist and patient versions (). Furthermore, the working alliance scores for all four scales can significantly detect individuals with suicidality (). There are also variations among the working alliance scores of each clinical conditions (Tables S5, S6, S7, S8, S9, and S10 in the Appendix, for statistical differences of the working alliance scores among conditions).
Temporal Dynamics of Working Alliance. We examine the temporal dynamics of the working alliance by mapping the trajectories in the alliance space for the three major scales (task, bond, and goal). Figures 3 and 4 illustrate the average trajectories of different psychiatric conditions. We observe that the trajectories of individuals with suicidality are more spread out in the bond and task scales, indicating significant discrepancy. This analysis provides a preliminary understanding of the temporal dynamics of the working alliance in different conditions, which can help therapists gain insights into the therapeutic process and guide further analysis.
3.3 Advantage of Inferred Psychometrics on the Diagnosis of Clinical Conditions
Next, we evaluate the usefulness of features derived from working alliance and dialogue turns in classifying psychotherapy sessions into four clinical labels. We employ two classifier backbones: Transformers [24] and Long Short-Term Memory Networks (LSTM) [25]. We compare three types of features: inferred scores of working alliance, pretrained document embedding, and both. Additionally, we explore the performance using dialogue turns from patients, therapists, and both patients and therapists.
@@ 3.3 Advantage of Inferred Psychometrics on the Diagnosis of Clinical Conditions
We further examine the influence of different features and embeddings on classification performance. The combined feature of inferred scores of pretrained document embedding and working alliance consistently outperforms the other feature types in both Transformer and LSTM models. When using SentenceBERT as the embedding, there is a modest benefit from training on patient turns alone, suggesting potential feature interference between therapist and patient working alliance information. The Transformers utilizing the combined feature of inferred scores and pretrained embedding, especially when using Doc2Vec as the sentence embedding, achieve the best performance. These results indicate the potential usefulness of inferred therapeutic or psychological state scores in downstream tasks, such as diagnosing clinical conditions. Future investigations could explore other downstream tasks and leverage the attention mechanism of Transformer blocks for interpretations.
3.4 Interpretable Clinical Insights Provided by Topic Modeling
@@
The evaluation metrics include topic embedding coherence (, , , ) and other measures such as coherence based on the asymmetrical confirmation measure and topic diversity. The results show variability in the rankings across the coherence metrics (Table S3 and S4 in the Appendix), but certain models consistently demonstrate higher topic coherence and diversity (for both metrics, the higher, the better). Following [28, 27], the topic diversity is computed as the proportion of unique words (PUW) with 1.0 being perfectly diverse, and the topic coherence is computed as a integrated log ratio between the co-document frequency and document frequency with higher being better (as correlated by expert ratings). Notably, the Wasserstein-based Topic Models and Embedded Topic Models exhibit relatively high coherence and diversity in these two metrics.
We further investigate the temporal dynamics of the learned Embedded Topic Model, which yields the best performance, on the text corpus of the entire psychotherapy dataset. By computing topic scores for each turn, we can analyze the dialogue dynamics within the topic space. Figure 4 presents the cumulative discrepancy of scores in the three scales (panel A) and of projections on the three principals topics (panel B) between patient-therapist dialogue pairs across topics. We observe in particular the largest growing discrepancy for suicidality, similarities between anxiety and depression, and a complementary behavior for schizophrenia, better represented by the trajectory in principal topic 1.
@@ 3.5 Major Themes in Psychotherapeutic Topics and Interventions
| Topics | Interventions | Examples |
Topic 0: Encouraging Change and Self-Reflection |
• Explore the potential for the individual to reconsider their viewpoints or habits. • Suggest maintaining certain behaviors for a period of time. • Discuss personal growth and maintaining a sense of self-awareness. |
“Well, I mean well keep doing the exercise.” |
Topic 1: Making a Case and Seeking Understanding |
• Discuss the importance of presenting a clear argument or perspective. • Inquire about interactions with a case manager or authority figure. • Explore the depth of personal experiences and the case’s significance in the person’s life. |
“You’ve made your case that weight’s important.” |
Topic 2: Maintaining Boundaries and Professionalism |
• Discuss the importance of establishing emotional boundaries. • Encourage the person to continue engaging in therapeutic exercises. • Reflect on maintaining professionalism and boundaries within the therapeutic relationship. |
“Yes, you keep it very professional; patient-doctor.” |
Topic 3: Discussing Coping Mechanisms and Playfulness |
• Encourage the person to engage in activities that bring them joy. • Explore social interactions and relationships, such as playdates and friendships. • Discuss using adaptive strategies like “playing it by ear” to navigate life’s uncertainties. |
“Play it up.” |
Topic 4: Focus on Specific Numbers and Memories |
• Discuss specific instances involving numbers (e.g., taking medication). • Explore memories associated with certain days or events. • Engage in conversation related to quantifiable aspects of the person’s life. |
“Just for like 1 day? Okay, how do you remember that day?” |
Topic 5: Exploring Anger and Emotions |
• Inquire about feelings of anger towards specific individuals. • Validate and explore the intensity of anger towards others. • Investigate triggers, circumstances, and potential underlying emotions contributing to anger. |
“Are you angry with him at all?” |
Topic 6: Discussing Drinking Habits |
• Inquire about the individual’s alcohol consumption. • Explore preferences related to different beverages, like soda and coffee. • Address drinking habits and patterns to gain insights into lifestyle choices. |
“Do you drink any, or almost none?” |
Topic 7: Philosophical Reflections and Communication |
• Engage in discussions about human nature and existence. • Reflect on personal experiences and projects. • Explore communication dynamics, including misperceptions and honesty. |
“Every human being?” |
Topic 8: Offering Hope and Support |
• Express empathy and hope for positive outcomes. • Acknowledge the individual’s progress and insights. • Provide encouragement to continue the journey of self-discovery and growth. |
“I hope so too. Take care.” |
Topic 9: Addressing Fatigue |
• Inquire about the reasons behind the individual feeling tired. • Explore the decision-making process that led to seeking therapy. • Elicit the motivations and factors that brought the person to the therapy session. |
“Tired a lot?” |
@@
In summary, the findings from our topic modeling analysis reveal several major themes that underpin the diverse array of therapeutic interactions observed in psychotherapy sessions. These themes provide a comprehensive framework for understanding the nuanced ways therapists guide clients through self-exploration, emotional processing, communication enhancement, and personal growth journeys.
@@ 3.6 Principal Topic Space in Psychotherapy Sessions
In the sessions with schizophrenia patients, and therapists in this dataset may prioritize addressing symptoms related to psychosis, managing hallucinations or delusions, and enhancing reality testing and medication adherence. The interventions attributed to therapists may involve psychoeducation about the illness, cognitive-behavioral interventions for managing symptoms, and collaborating with other healthcare providers.
@@
Insights for Clinicians. To explore the informative value of topics for therapeutic insights, we combine topic modeling with the inferred working alliance (Figure 5). By filtering the therapist turns with high topic scores, we plot the average working alliance scores for corresponding patient turns. We observe distinctions among the effects on patients’ working alliance across different topics and clinical conditions. For example, discussing tiredness and decision-making positively influences the bond and task scales in schizophrenia patients but has less impact on other patients. Additionally, discussing sickness, self-injuries, and coping mechanisms positively affects the task scale in depression patients and the goal scale in suicidal patients.
@@
If the clinicians discuss about principal component topic 1, “Emotional States and Mental Health”, it increases the TASK and BOND scales for depression patients, but decreases them for suicidal patients.
@@
If the clinicians discuss about principal component topic 2, “Personal Experiences and Relationships”, it increases GOAL scale for all patients, helps with the BOND scale for schizophrenia patients, but decreases the TASK and BOND scales for depression and suicidal patients.
@@
If the clinicians discuss about principal component topic 3, “Decision-Making, Self-Reflection and Personal Growth”, it might increase the BOND scale for depression patients and GOAL scale for schizophrenia patients, but decrease the BOND and TASK scales for suicidal patents.

### v2

@@ 2.7 Analyzing the Temporal Dynamics of Topics
To analyze the temporal dynamics of topics, we compute topic scores at the turn-level. We utilize the Embedded Topic Model (ETM) for this analysis, as it models each word with a categorical distribution based on the inner product between a word embedding and the embedding of its assigned topic [37]. We use the same Word2Vec word embeddings to embed both the topics and the dialogue turns, to then compute the cosine similarity between the embedded topic vector and the embedded turn vector. By applying these methods, which we term Temporal Topic Modeling (TMM), we obtain turn-resolution topic scores that capture the temporal dynamics of the topics discussed during the therapy session. These turn-level topic scores allow us to track the changes in topic relevance over time, providing insights into the progression of the therapy, the emergence of specific topics, and shifts in the focus of the conversation.
@@ 3.1 Insights on Patient-Doctor Relationship from Working Alliance Analysis
Patient-Therapist Consistency of Working Alliance. We investigate the consistency of the working alliance estimation between patients and therapists. Comparing the estimates, we observe that therapists tend to overestimate the working alliance overall. Specifically, therapists tend to overestimate the task and bond scales, but underestimate the goal scale. These differences are statistically significant (). We also find that the working alliance scores differ significantly between certain pairs of psychiatric conditions, such as anxiety and depression, and anxiety and schizophrenia, in both the therapist and patient versions (). Furthermore, the working alliance scores for all four scales can significantly detect individuals with suicidality (). There are also variations among the working alliance scores of each clinical conditions (Tables S4, S5, S6, S7, S8, and S9 in the Supplementary Materials, for statistical differences of the working alliance scores among conditions).
In TASK scale, significant differences were observed between anxiety and depression (p < 1.550e-26), between anxiety and schizophrenia (p < 3.835e-05), and between depression and suicidal tendencies (p < 1.900e-02) in the therapist turns. Similarly, patient turns revealed significant differences across all pairs of conditions, including between anxiety and depression (p < 1.796e-25), and schizophrenia and suicidal tendencies (p < 5.465e-14). These results suggest that therapists may approach task alignment differently depending on the patient’s psychiatric condition. The significant differences observed in both therapist and patient turns imply that therapists might align tasks more effectively with anxious and depressed patients compared to those with schizophrenia or suicidal tendencies.
In BOND scale, The therapist turns showed significant differences in the BOND scale between anxiety and depression (p < 1.600e-22), as well as between anxiety and schizophrenia (p < 2.998e-05). Patient turns revealed significant differences across almost all pairs, including between schizophrenia and depression (p < 3.883e-03), and schizophrenia and suicidal tendencies (p < 3.629e-34). These results suggest that therapists may perceive stronger bonds with anxious and depressed patients compared to those with schizophrenia or suicidal tendencies.
In GOAL scale, significant differences were found across all therapist pairs except between depression and schizophrenia (p = 4.959e-01). In the patient data, significant differences were found only between schizophrenia and anxiety (p < 8.006e-04), and between schizophrenia and depression (p < 4.519e-05). This suggests notable discrepancies in how therapists and patients perceive the alignment of treatment goals across these psychiatric conditions.
These variations are important because they reveal distinct patterns in the therapeutic alliance based on the clinical condition being treated. For example, patients with suicidal tendencies exhibit significantly larger discrepancies in their working alliance scores compared to other conditions, particularly in the TASK and BOND scales, but not as much from the therapist’s perspective. However, in the GOAL scale, therapists show significant differences across conditions, while patients do not perceive these differences as strongly. This mismatch may indicate that therapists are more attuned to differences in goal alignment across conditions than patients are, especially when it comes to conditions like anxiety and suicidality. Such discrepancies highlight the challenges therapists face in aligning therapeutic goals with certain patient populations, particularly those with suicidal tendencies, and suggest that more nuanced strategies are needed to ensure that both parties are on the same page regarding treatment goals.
Temporal Dynamics of Working Alliance. We examine the temporal dynamics of the working alliance by mapping the trajectories in the alliance space for the three major scales (task, bond, and goal). Specifically, we calculate the difference between patient and therapist scores to explore perceived misalignments in these aspects. Figures 3 and 4 illustrate the average trajectories across different psychiatric conditions, revealing several distinctive patterns.
For suicidality, the trajectory in bond and task scales shows a continuous decrease, moving increasingly negative over time, which suggests a growing misalignment as therapy progresses. This negative trend may reflect a disconnect in emotional rapport and task engagement between patient and therapist for this group. In contrast, the goal scale trajectory for suicidality shifts from zero to positive, implying an improving alignment on long-term goals. This may also reflect the particular challenges and variability in maintaining a strong alliance with suicidal patients, where initial engagement might wane over time as treatment progresses. However, it’s worth noting that our limited sample size for suicidality could introduce bias, so these patterns should be interpreted cautiously.
In the cases of anxiety and depression, we observe that both conditions start with a negative trajectory in the bond scale, which intensifies in the first half of the sessions. However, both conditions show a partial bounce-back in bond alignment later, even though the values remain predominantly negative, indicating a persistent gap in perceived emotional connection. For the task scale, the trajectories diverge: depression sessions increase from zero to positive, indicating improving alignment in task engagement, whereas anxiety sessions drop from zero to negative, suggesting a growing misalignment over time. This divergence could imply that task-focused interventions may need to be adapted differently for these two conditions over time.
For schizophrenia, the trajectories in both bond and goal scales display a consistent increase, remaining positive throughout the sessions. This suggests that for schizophrenia, patient-therapist alignment improves over time, particularly in terms of emotional connection and goal setting, which could indicate a progressive establishment of rapport and shared understanding in this population.
Clinical Implications and Roadmap. These distinct trajectories offer a preliminary roadmap for clinicians. By identifying prototypical alliance trajectories for each condition, therapists can begin to map individual sessions onto these broader patterns. For example, if a session with a suicidal patient diverges significantly from the typical trajectory (e.g., a sharper decline in the bond scale), this might signal a need for immediate intervention.
These patterns of alignment and misalignment provide clinicians with a roadmap for interpreting patient-therapist dynamics over time. Furthermore, therapists could use these trajectories to inform treatment adjustments, with specific attention to each scale:
In BOND scale, a downward trend in bond from the patient’s perspective, especially with suicidal or anxious patients, could prompt a focus on strengthening rapport and emotional connection. For suicidality, the continuous decline in bond alignment suggests an increasing emotional disconnect. Clinicians might focus on interventions that specifically aim to build rapport with suicidal patients, addressing the emotional aspects of therapy to prevent further drift.
In TASK scale, the divergence in task alignment between anxiety and depression implies that therapists might need to adapt task-related strategies differently for these conditions. The decreasing task alignment for anxiety, moving towards negative values, suggests re-evaluating the relevance of tasks to improve engagement, while the positive trend in depression could indicate opportunities to reinforce shared tasks.
In GOAL scale, the increasing goal alignment for schizophrenia, which remains positive, suggests that long-term goal alignment is progressively achieved, potentially validating goal-oriented therapeutic approaches. Conversely, the improving goal alignment in suicidality may signal a rare but valuable area of common ground between patient and therapist, which could be leveraged to foster hope and engagement.
Mapping individual sessions onto these trajectories could allow therapists to identify early warning signs of divergence from these patterns. For instance, observing a sudden drop in task or bond alignment could prompt a re-evaluation of the therapeutic approach for patients with similar conditions. Future work could focus on developing tools to track these alignment differences in real-time, enabling clinicians to make dynamic adjustments based on evolving alliance data.
@@
3.2 Inference of the Diagnosis of Clinical Conditions
Next, we evaluate the usefulness of features derived from working alliance and dialogue turns in classifying psychotherapy sessions into four clinical labels. This evaluation is intended solely for validation purposes and not as a diagnostic tool. Given that clinical diagnoses were already established and influenced the content of these transcripts, our goal is not to replace diagnostic standards like the Diagnostic and Statistical Manual of Mental Disorders (DSM), but rather to demonstrate that inferred alliance scores can correlate meaningfully with known clinical outcomes. This serves as a way to assess the predictive power of working alliance trajectories within the sessions, suggesting that the inferred scores can provide valuable insights into the types of patients being treated. These findings highlight the potential for working alliance measures to inform and guide treatment strategies based on patterns observable early in therapy sessions.
We employ two classifier backbones: Transformers [34] and Long Short-Term Memory Networks (LSTM) [35]. We compare three types of features: inferred scores of working alliance, pretrained document embedding, and both. Additionally, we explore the performance using dialogue turns from patients, therapists, and both patients and therapists.
@@ 3.2 Inference of the Diagnosis of Clinical Conditions
| Topics | Interventions | Examples |
| Topic 0: Encouraging Change and Self-Reflection | • Explore the potential for the individual to reconsider their viewpoints or habits. • Suggest maintaining certain behaviors for a period of time. • Discuss personal growth and maintaining a sense of self-awareness. | “Well, I mean well keep doing the exercise.” |
| Topic 1: Making a Case and Seeking Understanding | • Discuss the importance of presenting a clear argument or perspective. • Inquire about interactions with a case manager or authority figure. • Explore the depth of personal experiences and the case’s significance in the person’s life. | “You’ve made your case that weight’s important.” |
| Topic 2: Maintaining Boundaries and Professionalism | • Discuss the importance of establishing emotional boundaries. • Encourage the person to continue engaging in therapeutic exercises. • Reflect on maintaining professionalism and boundaries within the therapeutic relationship. | “Yes, you keep it very professional; patient-doctor.” |
| Topic 3: Discussing Coping Mechanisms and Playfulness | • Encourage the person to engage in activities that bring them joy. • Explore social interactions and relationships, such as playdates and friendships. • Discuss using adaptive strategies like “playing it by ear” to navigate life’s uncertainties. | “Play it up.” |
| Topic 4: Focus on Specific Numbers and Memories | • Discuss specific instances involving numbers (e.g., taking medication). • Explore memories associated with certain days or events. • Engage in conversation related to quantifiable aspects of the person’s life. | “Just for like 1 day? Okay, how do you remember that day?” |
| Topic 5: Exploring Anger and Emotions | • Inquire about feelings of anger towards specific individuals. • Validate and explore the intensity of anger towards others. • Investigate triggers, circumstances, and potential underlying emotions contributing to anger. | “Are you angry with him at all?” |
| Topic 6: Discussing Drinking Habits | • Inquire about the individual’s alcohol consumption. • Explore preferences related to different beverages, like soda and coffee. • Address drinking habits and patterns to gain insights into lifestyle choices. | “Do you drink any, or almost none?” |
| Topic 7: Philosophical Reflections and Communication | • Engage in discussions about human nature and existence. • Reflect on personal experiences and projects. • Explore communication dynamics, including misperceptions and honesty. | “Every human being?” |
| Topic 8: Offering Hope and Support | • Express empathy and hope for positive outcomes. • Acknowledge the individual’s progress and insights. • Provide encouragement to continue the journey of self-discovery and growth. | “I hope so too. Take care.” |
| Topic 9: Addressing Fatigue | • Inquire about the reasons behind the individual feeling tired. • Explore the decision-making process that led to seeking therapy. • Elicit the motivations and factors that brought the person to the therapy session. | “Tired a lot?” |
We further examine the influence of different features and embeddings on classification performance. The combined feature of inferred scores of pretrained document embedding and working alliance consistently outperforms the other feature types in both Transformer and LSTM models. When using SentenceBERT for sentence embeddings, we observed a modest improvement in classification performance when training on patient turns alone, as opposed to using both patient and therapist data together. This suggests that combining therapist and patient turns may introduce overlapping features that could interfere with the clarity of the inferred alliance scores. For clinicians, this finding highlights the potential value in separately analyzing patient dialogue, as it might yield more distinct and interpretable signals regarding the working alliance. The Transformers utilizing the combined feature of inferred scores and pretrained embedding, especially when using Doc2Vec as the sentence embedding, achieve the best performance.
These results indicate the potential usefulness of inferred therapeutic or psychological state scores in downstream tasks, such as supporting treatment planning and personalization. While this part of the study does not aim to replace clinical diagnosis, the inferred scores could offer additional predictive insights that complement existing clinical frameworks. For instance, in a transdiagnostic setting, these early session indicators might help clinicians understand which types of treatment plans are most suitable, even before a clear diagnosis is fully established. Future investigations could explore other downstream tasks and leverage the attention mechanism of Transformer blocks for further interpretation.
3.3 Topic Modeling Approach to Provide Interpretable Clinical Insights
@@
The evaluation metrics include topic embedding coherence (, , , ) and other measures such as coherence based on the asymmetrical confirmation measure and topic diversity. The results show variability in the rankings across the coherence metrics (Table S2 and S3 in the Supplementary Materials), but certain models consistently demonstrate higher topic coherence and diversity (for both metrics, the higher, the better). Following [38, 37], the topic diversity is computed as the proportion of unique words (PUW) with 1.0 being perfectly diverse, and the topic coherence is computed as an integrated log ratio between the co-document frequency and document frequency with higher being better (as correlated by expert ratings). Notably, the Wasserstein-based Topic Models and Embedded Topic Models exhibit relatively high coherence and diversity in these two metrics.
@@ 3.4 Major Themes in Psychotherapeutic Topics and Interventions
The major themes identified through this topic modeling approach represent the key therapeutic strategies and intervention techniques utilized by the therapist during sessions. These themes, directly related to intervention strategies such as cognitive reframing, emotional validation, or goal-setting, reflect the therapist’s active role in steering the therapeutic process. By focusing on the therapist’s contributions, we can better understand not just the content of the conversations, but also the clinical strategies that are emphasized in practice. These themes are particularly useful for extracting actionable insights that clinicians might find helpful in refining their approaches to treatment.
The findings from our topic modeling analysis reveal several major themes that underpin the diverse array of therapeutic interactions observed in psychotherapy sessions. These themes provide a comprehensive framework for understanding the nuanced ways therapists guide clients through self-exploration, emotional processing, communication enhancement, and personal growth journeys.
In contrast, the principal component topics, which are discussed in our next section, provide a broader view of the full dialogue between patient and therapist. These topics are learned from both therapist and patient turns and offer a general understanding of what is being talked about throughout the sessions, serving as a comprehensive reference for the entire discourse. Together, the major themes and principal component topics complement one another: the themes give us insight into the specific therapeutic techniques applied, while the topics provide a broader contextual framework for the overall content of therapy.
@@ 3.5 Principal Topic Space in Psychotherapy Sessions
In the sessions with schizophrenia patients, therapists in this dataset may prioritize addressing symptoms related to psychosis, managing hallucinations or delusions, and enhancing reality testing and medication adherence. The interventions attributed to therapists may involve psychoeducation about the illness, cognitive-behavioral interventions for managing symptoms, and collaborating with other healthcare providers.
@@
Temporal Dynamics of Therapy Topics. We further investigate the temporal dynamics of the learned Embedded Topic Model, which yields the best performance, on the text corpus of the entire psychotherapy dataset. By computing topic scores for each turn, we can analyze the dialogue dynamics within the topic space. Figure 4 presents the cumulative discrepancy of scores in the three scales (panel A) and of projections on the three principal topics (panel B) between patient-therapist dialogue pairs across topics.
Our analysis reveals several distinctive patterns. Specifically, we observe the largest growing discrepancy for suicidality, similarities in topic divergence between anxiety and depression, and a complementary pattern for schizophrenia, which is most prominent in principal topic 1 (PT1) related to emotional states and mental health.
For anxiety and depression, trajectories are more separable, with both conditions showing a consistent decline from zero to negative values in PT1 (emotional states and mental health), suggesting an increasing divergence between patients and therapists on these issues. In PT2 (personal experiences and relationships), the discrepancy also becomes gradually more negative, while in PT3 (decision-making and personal growth), both conditions show a shift from zero to positive. This indicates growing alignment in discussing personal growth, despite increasing misalignment on emotional and relational topics.
For schizophrenia, we observe a complementary pattern where PT2 shows an initial decline from zero to negative but stabilizes midway, suggesting early divergence in personal relationships. Meanwhile, PT1 shifts positively, starting after PT2 stabilizes, creating a twist in the trajectory. Additionally, PT3 initially dips negative before rebounding to positive. This evolving pattern implies that for schizophrenia, therapists and patients might initially diverge on personal experiences, but increasingly align on emotional health and self-reflection as sessions progress.
Suicidality shows the widest range of topic divergence. The trajectory for PT2 begins with a drop in alignment, indicating early divergence in discussions about personal experiences. PT3, meanwhile, initially rises from zero to positive but then sharply reverses to negative, suggesting fluctuating alignment on decision-making and personal growth topics. This spread-out pattern may reflect the more dynamic and complex nature of therapeutic engagement with suicidal patients, which can involve rapid shifts in focus.
We also observe that trajectories are generally more separable in anxiety and schizophrenia sessions, while depression sessions tend to be more entangled across topics. This separability may indicate distinct thematic focus patterns, which could offer therapists insights into how patients engage with different therapeutic areas based on their conditions.
Clinical Implications and Roadmap. By visualizing these trajectories, therapists can retrospectively examine session dynamics, providing an automatic annotation and notetaking mechanism. For junior therapists, this can serve as a valuable tool for skills development, allowing them to understand patient-therapist interactions and adjust strategies based on alignment patterns. Moreover, these insights could guide therapists in real-time, particularly when deviations from expected trajectories are observed. For example:
In PT1, the increasing divergence for anxiety and depression suggests that ongoing attention to emotional topics may be necessary to realign patient-therapist perceptions.
In PT2, early divergence in schizophrenia and suicidality indicates that a focus on personal relationships could be essential for establishing common ground.
In PT3, the positive alignment trajectory in schizophrenia and the erratic patterns in suicidality suggest that therapists might need to dynamically adjust strategies around personal growth and decision-making.
Future applications could involve real-time tracking of these discrepancies, enabling therapists to monitor and respond to evolving alignment patterns within sessions. This dynamic feedback could improve treatment outcomes by helping clinicians address emerging alignment gaps early in the therapeutic process.
@@
Topic-Alliance Associations as Actionable Insights for Clinicians. To explore the informative value of topics for therapeutic insights, we combine topic modeling with the inferred working alliance (Figure 5). By filtering the therapist turns with high topic scores, we plot the average working alliance scores for corresponding patient turns. We observe distinctions among the effects on patients’ working alliance across different topics and clinical conditions. For example, discussing tiredness and decision-making positively influences the bond and task scales in schizophrenia patients but has less impact on other patients. Additionally, discussing sickness, self-injuries, and coping mechanisms positively affects the task scale in depression patients and the goal scale in suicidal patients.
@@
If the clinicians discuss principal component topic 1, “Emotional States and Mental Health”, it increases the TASK and BOND scales for depression patients, but decreases them for suicidal patients.
@@
If the clinicians discuss principal component topic 2, “Personal Experiences and Relationships”, it increases GOAL scale for all patients, helps with the BOND scale for schizophrenia patients, but decreases the TASK and BOND scales for depression and suicidal patients.
@@
If the clinicians discuss principal component topic 3, “Decision-Making, Self-Reflection and Personal Growth”, it might increase the BOND scale for depression patients and GOAL scale for schizophrenia patients, but decrease the BOND and TASK scales for suicidal patents.